OpenStudy (anonymous):

a friend will collect ceramic and stuffed dogs. she has 248 presently. she wishes to add 8 a week to her collection until she reaches a total of 400 dogs. how long will it take for her to reach 400 dogs?

OpenStudy (anonymous):

Well the difference between 400 and 248 is 152. (400-248=152) Divide 152 by 8 to find how many weeks it will take to get to 400.

OpenStudy (anonymous):

19

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

write the following equation in slope intercept form and identify the slope and y intercept 2x+4y=7

OpenStudy (anonymous):

Change the equation to the form y=mx+c. To find the intercept on the y-axis let x=0 and sub it into your equation.